# Massive Fields Affected by Echoes: New Physics vs. Astrophysical Environment

### v1

@@ III Gravitational perturbations of squashed black holes in the Kaluza-Klein theory
Squashed Kaluza-Klein (KK) black holes Ishihara and Matsuno (2006) exhibit significant differences from standard Schwarzschild black holes, which are characterized by asymptotic flatness, and from black strings, even at energy levels where KK modes have not yet been activated. This makes squashed KK black holes a valuable model for exploring the dynamics of higher-dimensional spacetimes. Unlike their more conventional counterparts, these squashed black holes offer unique insights into higher dimensions, providing a glimpse into physics beyond the familiar four-dimensional framework.
One particularly noteworthy aspect of squashed KK black holes is their apparent stability Ishihara et al. (2008). This distinguishes them from other higher-dimensional solutions, such as black strings, which are prone to the Gregory-Laflamme instability Gregory and Laflamme (1993)—a phenomenon that can cause certain configurations to become dynamically unstable and fragment. The inherent stability of squashed KK black holes allows us to bypass this instability, making them more viable candidates.
The metric of the uncharged non-rotating squashed Kaluza-Klein black hole is Ishihara et al. (2008)
$ds^{2}$ $=$ $-F(\rho)dt^{2}+\frac{G(\rho)^{2}}{F(\rho)}d\rho^{2}$
$+\rho^{2}G(\rho)^{2}(\sigma_{1}^{2}+\sigma_{2}^{2})+\frac{\rho_{\infty}^{2}}{4G(\rho)^{2}}(\sigma_{3})^{2},$
$\sigma_{1}$ $=$ $-\sin{\psi}d\theta+\cos{\psi}\sin{\theta}d\phi\ ,$
$\sigma_{2}$ $=$ $\cos{\psi}d\theta+\sin{\psi}\sin{\theta}d\phi\ ,$ (10)
$\sigma_{3}$ $=$ $d\psi+\cos{\theta}d\phi\ ,$
and
$\rho_{\infty}^{2}=4\rho_{0}(\rho_{+}+\rho_{0})\ ,$
As shown in Kimura et al. (2008), the linearized perturbation equations of the gravitational field for zero modes with respect to the $SU(2)$ group are also reduced to the wave-like form with the following effective potentials,
$+64\rho^{3}(17\rho_{+}+2\rho)\rho_{0}+48\rho^{2}(32\rho_{+}+11\rho)\rho_{0}^{2}$
$\qquad+60\rho(13\rho_{+}+12\rho)\rho_{0}^{3}+9(9\rho_{+}+35\rho)\rho_{0}^{4}\Big{]}\ ,$
$V_{1}$ $=$ $\left(1-\frac{\rho_{+}}{\rho}\right)\bigg{[}\frac{1}{\rho_{+}\rho_{0}}+\frac{7(-\rho_{+}+\rho)}{16(\rho+\rho_{0})^{3}}$
$+\frac{17\rho_{+}-19\rho}{8\rho(\rho+\rho_{0})^{2}}+\frac{9(-3\rho_{+}+7\rho)}{16\rho^{2}(\rho+\rho_{0})}+\frac{\rho_{+}-\rho}{\rho_{+}^{2}\rho+\rho_{+}\rho\rho_{0}}$
$-\frac{8(\rho_{+}-\rho)^{2}\rho}{(\rho_{+}-2\rho)(\rho_{+}\rho_{0}-\rho(\rho+2\rho_{0}))^{2}}$
$-\frac{2(\rho_{+}+2\rho)}{(\rho_{+}-2\rho)(\rho_{+}\rho_{0}-\rho(\rho+2\rho_{0}))}\bigg{]}\ ,$
$V_{2}$ $=$ $\frac{-(\rho_{+}-\rho)}{16\rho^{3}\rho_{0}(\rho_{+}+\rho_{0})(\rho+\rho_{0})^{3}}\Big{[}64\rho^{5}$
$+256\rho^{4}\rho_{0}-32\rho^{3}\rho_{0}(\rho_{+}-11\rho_{0})+9\rho_{+}\rho_{0}^{3}(\rho_{+}+\rho_{0})$
$+8\rho^{2}\rho_{0}(2\rho_{+}^{2}-5\rho_{+}\rho_{0}+25\rho_{0}^{2})$
$+\rho\rho_{0}^{2}(20\rho_{+}^{2}-9\rho_{+}\rho_{0}+35\rho_{0}^{2})\Big{]}\ .$
From Figures 4 and 5, we observe that, similar to the case of massive scalar field perturbations, when the bump is located at a considerable distance from the black hole, the echoes start to modify the power-law behavior of the asymptotic tails, turning them into an oscillatory pattern. As the bump moves closer to the event horizon, the echoes appear earlier and begin to influence the ringing phase. Similar to the case of the Schwarzschild black hole, to observe echoes, the geometry needs to be modified at a distance from the radiation zone (Fig. 4). A potential bump of the same size affects the $K=1$ perturbation profile more significantly due to the smaller effective mass.
IV Schwarzschild-like wormholes
@@
Figure 6: Effective potential (left) and time-domain profile for a massive scalar field perturbations around the Schwarzschild black hole (blue): $r_{0}=1$, $\ell=1$, $\mu=1$ and the wormhole with $r_{t}=1.001$. The observer is situated at $r_{p}=11.721$, and the center of the Gaussian wave-package at the throat of the wormhole, $p=20$. The central panel shows the initial stage of the signal while the right panel shows also the asymptotically late-time behavior.
In this work, we consider $Z_{2}$-symmetric Schwarzschild-like wormholes Cardoso et al. (2016b) by assuming a Schwarzschild spacetime everywhere except at the wormhole throat, where a thin shell of matter is placed. The throat is positioned near the event horizon, at $r_{t}=r_{0}+\epsilon$ and we examine the region $r>r_{t}$ on both sides of the throat. For massless fields, the effective potential in this configuration takes the form of a double-peak potential, where the single peak in the Schwarzschild spacetime is split into two due to the symmetry of the radial coordinate. The echoes in this case arise from secondary scattering and reflection between the two peaks.
@@ V Conclusions
Our work could be extended by including an analysis of the decay of massive fields in other scenarios that produce echoes, such as black hole/wormhole transitions in the brane-world model considered in Bronnikov and Konoplya (2020); Bolokhov and Konoplya (2024) or various black hole models immersed in the galactic halo Konoplya et al. (2024), either within specific models of the dark matter or an approximate solution for a black hole surrounded by a halo Konoplya and Zhidenko (2022).

### v2

@@ 3 Gravitational perturbations of squashed black holes in the Kaluza-Klein theory
From Figures 5 and 6, we observe that, similar to the case of massive scalar field perturbations, when the bump is located at a considerable distance from the black hole, the echoes start to modify the power-law behavior of the asymptotic tails, turning them into an oscillatory pattern. As the bump moves closer to the event horizon, the echoes appear earlier and begin to influence the ringing phase. Similar to the case of the Schwarzschild black hole, to observe echoes, the geometry needs to be modified at a distance from the radiation zone (Fig. 5). A potential bump of the same size affects the $K=1$ perturbation profile more significantly due to the smaller effective mass.
4 Schwarzschild-like wormholes
In this work, we consider $Z_{2}$-symmetric Schwarzschild-like wormholes [60] by assuming a Schwarzschild spacetime everywhere except at the wormhole throat, where a thin shell of matter is placed. The throat is positioned near the event horizon, at $r_{t}=r_{0}+\epsilon$ and we examine the region $r>r_{t}$ on both sides of the throat. For massless fields, the effective potential in this configuration takes the form of a double-peak potential, where the single peak in the Schwarzschild spacetime is split into two due to the symmetry of the radial coordinate. The echoes in this case arise from secondary scattering and reflection between the two peaks.
@@ 4 Schwarzschild-like wormholes
Figure 7: Effective potential (left) and time-domain profile for a massive scalar field perturbations around the Schwarzschild black hole (blue): $r_{0}=1$, $\ell=1$, $\mu=1$ and the wormhole with $r_{t}=1.001$ (red). The observer is situated at $r_{p}=11.721$, and the center of the Gaussian wave-package at the throat of the wormhole, $p=20$. The central panel shows the initial stage of the signal while the right panel shows also the asymptotically late-time behavior.
@@ 5 Conclusions
There is one remark regarding our use of the term “echoes”. In the massive case, when the mass is sufficiently large, the effective potential does not develop two distinct peaks. However, secondary scattering off the bump still occurs, leading to a late-time modulation of the signal, leading to the oscillatory power-law envelope. This behavior differs from the massless case, where the modulation appears as a sequence of “ringdowns” called echoes. We nevertheless adopt the term “echoes” here, as we are referring to modifications of the late-time signal caused by the presence of the bump.
Our work could be extended by including an analysis of the decay of massive fields in other scenarios that produce echoes, such as black hole/wormhole transitions in the brane-world model considered in [9, 67] or various black hole models immersed in the galactic halo [68], either within specific models of the dark matter or an approximate solution for a black hole surrounded by a halo [69].